# Hybrid quantum–classical matrix-product state and Lanczos methods for electron–phonon systems with strong electronic correlations: Application to disordered systems coupled to Einstein phonons

### v1

@@ I Introduction
Electron–phonon coupling is ubiquitous in many phenomena in solid-state physics and quantum chemistry. In solid-state physics, the importance of phonons for electronic properties is evident in transport measurements, optical excitations, for the formation of Cooper pairs in conventional BCS superconductors [BCS1957], and in the physics of polarons [Frohlich1954]. More recently, the role of phonons as a relaxation channel for electronic excitations became experimentally accessible due to the development of ultrafast dynamics techniques [Giannetti2016] and is also discussed in the context of energy-conversion processes [Pastor2022]. Subsequently, with similar tools, the possibility of modifying electronic properties via optically targeting specific phonon modes has been explored (see [Bloch2022] for a review and [Rini2007, Tobey2008] for examples). In quantum chemistry, the role of vibrations is obvious for the excitation spectra of molecules and for certain reaction processes, to name a few examples [Nibbering2005].
@@
The theoretical treatment of electron–phonon problems becomes particularly challenging when additional interactions or degrees of freedom need to be accounted for. For instance, one can think of materials with orbital, spin, and phonon degrees of freedom with a substantial mutual coupling [Dagotto2001, Tokura2006]. In addition, whenever bare or induced electronic interactions push the system out of the validity range of perturbation theory, mean-field or density-functional theory approaches, then, there is a limited set of unbiased many-body techniques.
@@
Consequently, there has been recent activity in combining computational many-body methods that can capture strong electronic correlations with an efficient treatment of the phonon sector. These include (quantum)-Monte-Carlo methods [Weber2021, Cohen-Stead2023], dynamical mean-field theory [Picano2023], hierarchical equations of motions [Jankovic2022], Lanczos techniques [Bonca1999, Dorfner2015], Green’s function methods [Goodvin2008], matrix-product state approaches [Jeckelmann1998, Guo2011, Brockt2015, Paeckel2017, Jansen2020, Stolpp2021, Jansen2022, Jansen2023].
@@
In our work, we are concerned with wave-function methods in finite-dimensional Hilbert spaces such as Lanczos [Manmana2005] or matrix-product state (MPS) methods [Schollwoeck2011]. In previous MPS and density-matrix renormalization group developments, several ways of dealing with quantum phonons were introduced, including the pseudosite technique [Jeckelmann1998], local basis optimization [Zhang1998], and projected purification [Paeckel2017]. However, the regime of slow phonons with phonon energies much smaller than the bandwidth remains hard to access for these approaches [Stolpp2021]. In typical solid-state settings, however, this adiabatic regime is often the physically relevant one [Giustino2017].
@@
As a compromise, one can resort to well-established semi-classical methods that are widely used in connection with ab initio approaches [Lively2023] or in quantum chemistry. The simplest such technique is the multi-trajectory Ehrenfest method (MTE) [Aleksandrov1981, Tully1998, Nielsen2001], which treats the phonon dynamics classically. The quantum mechanical uncertainty is accounted for by a sampling of the initial coordinates and momenta from the initial quantum state probability distributions (often assumed to be the ground states), resulting in independent trajectories. We present the successful implementation of the MTE approach with a full quantum mechanical treatment of strongly correlated electrons in both time-dependent Lanczos and MPS simulations. This complements the usage of MTE with electrons that are only weakly electronically correlated and where such correlations are dealt with approximations such as Hartree-Fock (see, e.g., [Li2005]) or density functional theory [Lively2023].
@@
MTE is formally equivalent to the truncated Wigner approximation [Polkovnikov2010] of the phonon sector combined with a mean-field treatment of the electron–phonon coupling. In an earlier work, it was shown that MTE for the one-dimensional Holstein model yields numerically accurate results for electronic observables in the adiabatic regime for certain initial conditions [TenBrink2022]. In that case, the actual problem solved with MTE consists of non-interacting electrons whose wave function depends on the time-dependent momenta and coordinates of the oscillators. Thus, essentially, single-electron orbitals are propagated. Here, we extend this to a many-body problem of electrons that interact already in the absence of an electron–phonon coupling and hence a propagation of the fully many-body wave function is required. This can be accomplished in one spatial dimension.
@@
A related algorithm has recently been presented by Manawadu et al. [Manawadu2022, Manawadu2023], where a time-dependent DMRG algorithm is combined with Ehrenfest dynamics for phonons, in a single-trajectory formulation, based on the original block-version of the density matrix renormalization group method [White1992, White1993]. Our implementation is based on MPS and therefore, can easily be adapted to new algorithmic developments that are largely formulated in the MPS language. In particular, one can incorporate modern time-evolution schemes such as the time-dependent variational principle [Paeckel2019].
@@
We note that similar ideas were recently pursued in Monte-Carlo simulations [Weber2021] and in dynamical mean-field theory [Picano2023], yet both lack an exact incorporation of electronic correlations. There are also many extensions to MTE that account for its known shortcomings such as surface-hopping methods [Subotnik2016] or coupled-trajectory methods such as the multi-configurational Ehrenfest approach [Shalashilin2009, Shalashilin2010]. While we consider it worth to pursue such extensions in the future, we here aim at establishing the simplest and computationally inexpensive approach first.
@@
As an application, we consider the temporal decay of charge density wave (CDW) order. Such CDW order can either be present in the equilibrium phase diagram of interacting electrons and may decay due to external perturbations or we may envisage the initial CDW as being a result of some prior perturbations, e.g., in a pump-and-probe experiment (see, e.g., [Huang2024]). For the case of interest here, i.e., one-dimensional systems, the decay of CDW order of spinless fermions due electronic interactions and/or electron–phonon interactions has been studied previously [Hashimoto2017, Stolpp2020].
@@
Moreover, we are specifically interested in the decay of CDW order in systems with quenched and uncorrelated disorder as this relates to the stability of a putative many-body localization (MBL) in a system of interacting electrons [Nandkishore2015, Abanin2019]. We will show clear evidence that the coupling to classical phonons in the adiabatic regime leads to a decay of CDW order, indicating the instability of MBL in such a system. Furthermore, we will study the spreading of a single polaron in a disorder potential (see [Bronold2002, Bronold2004, Das2008, Berciu2010, Berciu2012, Tozer2014, Kloss2019]) which suggests sub-diffusive dynamics, consistent with [Prelovsek2018]. Our analysis of the CDW order decay hints at the presence of subdiffusive dynamics even at a finite filling, at least at intermediate times.
@@
The broader topic of thermalization in electron–phonon systems, studied from a model Hamiltonian perspective, has, of course, a long history, beyond what can reasonably be reviewed in this article. Recently, for instance, the dynamics after quenches in the electron–phonon coupling were studied using DMFT [Murakami2015]. The addition of disorder into an electron–phonon coupled system, already in the absence of direct electronic interactions, leads to partial delocalization as evidenced in, e.g., [DiSante2017], related to bad metal physics and variable-range hopping [Fleishman1978]. Note that the stability of many-body localization in the presence of a variety of different baths has been studied previously [HuseNandkishore2015, Sarang2017, Bar2022, Abanin2022, SerbynAbanin2022, Sierant2023b].
@@ II Model
$\mathcal{O}_{\mathrm{CDW}}=\frac{1}{N}\sum\limits_{\ell=1}^{L}(-1)^{\ell}\langle\hat{n}_{\ell}\rangle\,$, (8)
@@
In the context of MBL, one often studies the so-called imbalance $I$ [Schreiber2015], defined as
$I_{\mathrm{CDW}}$ $=\frac{N_{\text{even}}-N_{\text{odd}}}{N_{\text{even}}+N_{\text{odd}}}\,$, (10)
$N_{\text{even}}$ $=\sum_{\ell=1}^{L/2}\langle\hat{n}_{2\ell}\rangle\,$, (11)
$N_{\text{odd}}$ $=\sum_{\ell=1}^{L/2}\langle\hat{n}_{2\ell-1}\rangle\,.$ (12)
@@ III Methods
In this section, we provide a brief summary of the MTE method in Sec. III.1. For a detailed discussion, see [TenBrink2022] for a review. We then describe the incorporation of MTE into time-dependent Lanczos [Manmana2005] and TEBD [Vidal2004] in Secs. III.2 and III.4, respectively. We discuss the convergence of both hybrid methods.
@@ III.1.1 Basic idea and algorithm
The multi-trajectory Ehrenfest method (MTE) can be derived as a special case of the quantum–classical Liouville (QCL) equation [Aleksandrov1981, Nielsen2001], with two essential approximations: The coupling between the electronic subsystem and the phonon environment is treated in a mean-field approximation and the phonon dynamics is treated classically [Grunwald2009]. Intuitively, the latter is a reasonable assumption when the phonon dynamics is much slower then the electronic dynamics. The QCL equation is based on treating the phonon environment in the phase space formalism of quantum mechanics and can be used to approximate the dynamics of the partially transformed Wigner function [Wigner1932, Imre1967]. The full derivation of MTE within this formalism is illuminating, yet for the purpose of our work, we will only need the key algorithmic steps and a discussion of control parameters and convergence. We refer the reader to [TenBrink2022] and references therein for a more detailed account and a comparison of MTE to other trajectory-based methods.
We first discuss the single-trajectory version [Tully1998] and provide the equations of motion for the Hamiltonian from Eq. 1 for the quantum evolution of the electronic state $|\psi(t)\rangle$ and the classical evolution of the set of coordinates and momenta $\{x_{\ell},p_{\ell}\}$ of the oscillators. The electronic state depends on the $\{x_{\ell},p_{\ell}\}$, that is, $|\psi(t)\rangle=|\psi_{\{x_{\ell},p_{\ell}\}}(t)\rangle$. The classical and quantum Hamiltonians are
@@
The MTE scheme consist of simulating many single trajectories for a set of given initial conditions $\{x_{\ell}(t=0),p_{\ell}(t=0)\}$. The initial conditions for each trajectory are drawn from the quantum mechanical Wigner function of the initial state in the phonon sector. Subsequently, observables are computed for individual trajectories and then averaged over sufficiently many trajectories, with $N_{\text{traj}}$ the number of trajectories. Since the equations of motion, by construction, aim to conserve the properties of the Wigner function, we expect that the trajectory of the observable will converge—within the aforementioned approximations—to the quantum operator expectation value with $\sim 1/\sqrt{N_{\mathrm{traj}}}$ (see the statistical convergence analysis in [TenBrink2022]). Concretely, for our initial state, we sample $\{x_{\ell},p_{\ell}\}$ from a multivariate Gaussian distribution
@@
The computational effort for systems with disorder, treated later in this manuscript, is considerable because one has to generate multiple trajectories for each disorder realizations. Indeed, in such cases, we need to trade-off between the desired accuracy and expended computational effort. We notice that for realizations with strong disorder, the fluctuations in the observables between trajectories of different initial states are small compared to the fluctuations between different disorder realizations. Therefore we choose to sample each disorder realization only with $N_{\mathrm{traj}}={200}$ trajectories with initial conditions drawn from Eq. 18 for each of $N_{\mathrm{dis}}=$2002000\text{\,}$\$ different disorder realizations. This allows us to reduce the absolute error in averaged observables due to insufficient sampling of the simulations to $\sim 10^{-2}$, which we estimate by resampling through bootstrapping.
@@ III.1.2 Time propagation
Obviously, there is a choice in how the quantum subsytem is propagated. Most applications of MTE treat the electrons as non-interacting or weakly interacting, using Hartree-Fock [Li2005] or ab-initio methods [Lively2023]. In our work we implement two time-propagation schemes that treat electronic correlations accurately on finite systems, Lanczos and TEBD.
@@ III.1.3 Discussion of strength and weaknesses of MTE
MTE is one of the simplest trajectory based methods [Ehrenfest1927, Tully1998, Kirrander2020], stable, and straightforward to implement. Some limitations can be understood in the Born-Oppenheimer picture. For instance, in MTE, electrons experience the average force from several Born-Oppenheimer surfaces and do not see the effects of the individual Born-Oppenheimer surfaces and thus cannot account for splitting of wave packets, etc.
Another known issue in independent trajectory methods is overcoherence, namely the approximations ignore physical processes that cause decoherence [Kaeb2002]. Due to the classical nature of the time-evolution in the environment a dephasing or interference of quantum modes can not happen, resulting in overcoherent dynamics [Kaeb2002]. Some of the more sophisticated trajectory-based methods are designed to overcome these shortcomings of MTE, see [TenBrink2022] and references therein, however, in many cases, these modifications are based on further assumptions about the system and/or can make the implementation of the algorithm significantly more complex and less scalable.
@@
Despite the inherent limitations, nevertheless, for the case of adiabatic phonons $\unit\omega_{0}\ll\unit{}$, there are examples where MTE agrees with exact results. For instance for initial product states in the computational basis, electronic observables exhibit very small errors when compared to full quantum methods, while phonon observables qualitatively follow the correct dynamics [TenBrink2022]. These observations refer to expectation values only. The influence of the initial state on the quality of the agreement between MTE and exact methods has recently been discussed by Paprotzki et al. [Paprotzki2023]. Certain terms in the short-time expansion cancel out for initial states that are diagonal in the computational basis, yet play a role for, e.g., Slater determinants in the electronic quasimomentum eigenbasis. For other initial states with a finite density of electrons, e.g., a Fermi sea, phononic occupations may even become negative at large values of $\unit\omega_{0}\sim 2\unit{}$ [Paprotzki2023]. This relates to the well-known fact that in the long-time, stationary state, occupations obtained from Ehrenfest dynamics may deviate from the correct result [Parandekar2005, Parandekar2006].
Concerning the mere convergence with respect to its actual numerical control parameters, the statistical and numerical errors controlled by $N_{\mathrm{traj}}$ and $\Delta t$, convergence with respect to these parameters can easily be achieved in MTE. This makes MTE a technically simple-to-implement and to control technique, rendering it a reasonable starting point for hybrid methods for quantum many-body problems. While beyond the scope of paper, we envision that in the future, Lanczos or TEBD may also be combined with coupled trajectory methods like multi-configurational Ehrenfest [Shalashilin2009, Shalashilin2010] which at least has internal control parameters that allow convergence to the physically correct results in principle, as was demonstrated in [TenBrink2022].

### v2

@@ I Introduction
Electron–phonon coupling is ubiquitous in many phenomena in solid-state physics and quantum chemistry. In solid-state physics, the importance of phonons for electronic properties is evident in transport measurements, optical excitations, for the formation of Cooper pairs in conventional BCS superconductors [4], and in the physics of polarons [27]. More recently, the role of phonons as a relaxation channel for electronic excitations became experimentally accessible due to the development of ultrafast dynamics techniques [28] and is also discussed in the context of energy-conversion processes [79]. Subsequently, with similar tools, the possibility of modifying electronic properties via optically targeting specific phonon modes has been explored (see [6] for a review and [83, 98] for examples). In quantum chemistry, the role of vibrations is obvious for the excitation spectra of molecules and for certain reaction processes, to name a few examples [72].
@@
The theoretical treatment of electron–phonon problems becomes particularly challenging when additional interactions or degrees of freedom need to be accounted for. For instance, one can think of materials with orbital, spin, and phonon degrees of freedom with a substantial mutual coupling [17, 99]. In addition, whenever bare or induced electronic interactions push the system out of the validity range of perturbation theory, mean-field or density-functional theory approaches, then, there is a limited set of unbiased many-body techniques.
@@
Consequently, there has been recent activity in combining computational many-body methods that can capture strong electronic correlations with an efficient treatment of the phonon sector. These include (quantum)-Monte-Carlo methods [103, 15], dynamical mean-field theory [80], hierarchical equations of motions [44], Lanczos techniques [7, 21], Green’s function methods [31], matrix-product state approaches [48, 34, 11, 74, 45, 95, 46, 47].
@@
In our work, we are concerned with wave-function methods in finite-dimensional Hilbert spaces such as Lanczos [64] or matrix-product state (MPS) methods [85]. In previous MPS and density-matrix renormalization group developments, several ways of dealing with quantum phonons were introduced, including the pseudosite technique [48], local basis optimization [108], and projected purification [74]. However, the regime of slow phonons with phonon energies much smaller than the bandwidth remains hard to access for these approaches [95]. In typical solid-state settings, however, this adiabatic regime is often the physically relevant one [29].
@@
As a compromise, one can resort to well-established semi-classical methods that are widely used in connection with ab initio approaches [58] or in quantum chemistry. The simplest such technique is the multi-trajectory Ehrenfest method (MTE) [2, 101, 73], which treats the phonon dynamics classically. The quantum mechanical uncertainty is accounted for by a sampling of the initial coordinates and momenta from the initial quantum state probability distributions (often assumed to be the ground states), resulting in independent trajectories. We present the successful implementation of the MTE approach with a full quantum mechanical treatment of strongly correlated electrons in both time-dependent Lanczos and MPS simulations. This complements the usage of MTE with electrons that are only weakly electronically correlated and where such correlations are accounted for with approximations such as Hartree-Fock (see, e.g., [57]) or density functional theory [58].
@@
MTE is formally equivalent to the truncated Wigner approximation [81] of the phonon sector combined with a mean-field treatment of the electron–phonon coupling. In an earlier work, it was shown that MTE for the one-dimensional Holstein model yields numerically accurate results for electronic observables in the adiabatic regime for certain initial conditions [97]. In that case, the actual problem solved with MTE consists of non-interacting electrons whose wave function depends on the time-dependent momenta and coordinates of the oscillators. Thus, essentially, single-electron orbitals are propagated. Here, we extend this to a many-body problem of electrons that interact already in the absence of an electron–phonon coupling and hence a propagation of the fully many-body wave function is required. This can be accomplished in one spatial dimension.
@@
A related algorithm has recently been presented by Manawadu et al. [63, 62], where a time-dependent DMRG algorithm is combined with Ehrenfest dynamics for phonons, in a single-trajectory formulation, based on the original block-version of the density matrix renormalization group method [104, 105]. Our implementation is based on MPS and therefore, can easily be adapted to new algorithmic developments that are largely formulated in the MPS language. In particular, one can incorporate modern time-evolution schemes such as the time-dependent variational principle [75].
@@
We note that similar ideas were recently pursued in Monte-Carlo simulations [103] and in dynamical mean-field theory [80], yet both lack an exact incorporation of electronic correlations. There are also many extensions to MTE that account for its known shortcomings such as surface-hopping methods [96] or coupled-trajectory methods such as the multi-configurational Ehrenfest approach [89, 90]. While we consider it worth to pursue such extensions in the future, we here aim at establishing the simplest and computationally inexpensive approach first.
@@
As an application, we consider the temporal decay of charge density wave (CDW) order. Such CDW order can either be present in the equilibrium phase diagram of interacting electrons and may decay due to external perturbations or we may envisage the initial CDW as being a result of some prior perturbations, e.g., in a pump-and-probe experiment (see, e.g., [38]). For the case of interest here, i.e., one-dimensional systems, the decay of CDW order of spinless fermions due electronic interactions and/or electron–phonon interactions has been studied previously [36, 94].
@@
Moreover, we are specifically interested in the decay of CDW order in systems with quenched and uncorrelated disorder as this relates to the stability of a putative many-body localization (MBL) in a system of interacting electrons [71, 1]. We will show clear evidence that the coupling to classical phonons in the adiabatic regime leads to a decay of CDW order, indicating the instability of MBL in such a system. Furthermore, we will study the spreading of a single polaron in a disorder potential (see [13, 12, 18, 5, 22, 100, 52]) which suggests sub-diffusive dynamics, consistent with [82]. Our analysis of the CDW order decay hints at the presence of subdiffusive dynamics even at a finite filling, at least at intermediate times.
@@
The broader topic of thermalization in electron–phonon systems, studied from a model Hamiltonian perspective, has, of course, a long history, beyond what can reasonably be reviewed in this article. Recently, for instance, the dynamics after quenches in the electron–phonon coupling were studied using DMFT [70]. The combination of disorder and electron–phonon coupled system, already in the absence of direct electronic interactions, leads to partial delocalization as evidenced in, e.g., [20], related to bad metal physics and variable-range hopping [26]. Note that the stability of many-body localization in the presence of a variety of different baths has been studied previously [39, 32, 56, 10, 9, 92].
@@ II Model
$\mathcal{O}_{\mathrm{CDW}}=\frac{1}{N}\sum\limits_{\ell=1}^{L}(-1)^{\ell{+1}}\langle\hat{n}_{\ell}\rangle\,$, (8)
@@
In the context of MBL, one often studies the so-called imbalance $I$ [87], defined as
$I_{\mathrm{CDW}}$ $={\frac{N_{\text{odd}}-N_{\text{even}}}{N_{\text{odd}}+N_{\text{even}}}}\,$, (10)
${N_{\text{odd}}}$ ${=\sum_{\ell=1}^{L/2}\langle\hat{n}_{2\ell-1}\rangle\,,}$ (11)
${N_{\text{even}}}$ ${=\sum_{\ell=1}^{L/2}\langle\hat{n}_{2\ell}\rangle\,.}$ (12)
@@ III Methods
In this section, we provide a brief summary of the MTE method in Sec. III.1. For a detailed discussion, see [97] for a review. We then describe the incorporation of MTE into time-dependent Lanczos [64] and TEBD [102] in Secs. III.2 and III.4, respectively. We discuss the convergence of both hybrid methods.
@@ III.1.1 Basic idea and algorithm
The multi-trajectory Ehrenfest method (MTE) can be derived as a special case of the quantum–classical Liouville (QCL) equation [2, 73], with two essential approximations: The coupling between the electronic subsystem and the phonon environment is treated in a mean-field approximation and the phonon dynamics is treated classically [33]. Intuitively, the latter is a reasonable assumption when the phonon dynamics is much slower then the electronic dynamics. The QCL equation is based on treating the phonon environment in the phase space formalism of quantum mechanics and can be used to approximate the dynamics of the partially transformed Wigner function [106, 42]. The full derivation of MTE within this formalism is illuminating, yet for the purpose of our work, we will only need the key algorithmic steps and a discussion of control parameters and convergence. We refer the reader to [97] and references therein for a more detailed account and a comparison of MTE to other trajectory-based methods.
We first discuss the single-trajectory version [101] and provide the equations of motion for the Hamiltonian from Eq. 1 for the quantum evolution of the electronic state $|\psi(t)\rangle$ and the classical evolution of the set of coordinates and momenta $\{x_{\ell},p_{\ell}\}$ of the oscillators. The electronic state depends on the $\{x_{\ell},p_{\ell}\}$, that is, $|\psi(t)\rangle=|\psi_{\{x_{\ell},p_{\ell}\}}(t)\rangle$. The classical and quantum Hamiltonians are
@@
The MTE scheme consist of simulating many single trajectories for a set of given initial conditions $\{x_{\ell}(t=0),p_{\ell}(t=0)\}$. The initial conditions for each trajectory are drawn from the quantum mechanical Wigner function of the initial state in the phonon sector. Subsequently, observables are computed for individual trajectories and then averaged over sufficiently many trajectories, with $N_{\text{traj}}$ the number of trajectories. Since the equations of motion, by construction, aim to conserve the properties of the Wigner function, we expect that the trajectory of the observable will converge—within the aforementioned approximations—to the quantum operator expectation value with $\sim 1/\sqrt{N_{\mathrm{traj}}}$ (see the statistical convergence analysis in [97]). Concretely, for our initial state, we sample $\{x_{\ell},p_{\ell}\}$ from a multivariate Gaussian distribution
@@
The computational effort for systems with disorder, treated later in this manuscript, is considerable because one has to generate multiple trajectories for each disorder realizations. Indeed, in such cases, we need to trade-off between the desired accuracy and expended computational effort. We notice that for realizations with strong disorder, the fluctuations in the observables between trajectories of different initial states are small compared to the fluctuations between different disorder realizations. Therefore we choose to sample each disorder realization only with $N_{\mathrm{traj}}={200}$ trajectories with initial conditions drawn from Eq. 18 for each of $N_{\mathrm{dis}}=$2002000\text{\,}$\$ different disorder realizations. This allows us to reduce the absolute error in $\mathcal{O}_{\mathrm{CDW}}$, due to insufficient sampling of trajectories, to $\sim 10^{-2}$, which we estimate by resampling through bootstrapping.
@@ III.1.2 Time propagation
Obviously, there is a choice in how the quantum subsytem is propagated. Most applications of MTE treat the electrons as non-interacting or weakly interacting, using Hartree-Fock [57] or ab-initio methods [58]. In our work we implement two time-propagation schemes that treat electronic correlations accurately on finite systems, Lanczos and TEBD.
@@ III.1.3 Discussion of strength and weaknesses of MTE
MTE is one of the simplest trajectory based methods [23, 101, 51], stable, and straightforward to implement. Some limitations can be understood in the Born-Oppenheimer picture. For instance, in MTE, electrons experience the average force from several Born-Oppenheimer surfaces and do not see the effects of the individual Born-Oppenheimer surfaces and thus cannot account for splitting of wave packets, etc.
Another known issue in independent trajectory methods is overcoherence, namely the approximations ignore physical processes that cause decoherence [50]. Due to the classical nature of the time-evolution in the environment a dephasing or interference of quantum modes can not happen, resulting in overcoherent dynamics [50]. Some of the more sophisticated trajectory-based methods are designed to overcome these shortcomings of MTE, see [97] and references therein. However, in many cases, these modifications are based on further assumptions about the system and/or can make the implementation of the algorithm significantly more complex and less scalable.
@@
Despite the inherent limitations, nevertheless, for the case of adiabatic phonons $\unit\omega_{0}\ll\unit{}$, there are examples where MTE agrees with exact results. For instance for initial product states in the computational basis, electronic observables exhibit very small errors when compared to full quantum methods, while phonon observables qualitatively follow the correct dynamics [97]. These observations refer to expectation values only. The influence of the initial state on the quality of the agreement between MTE and exact methods has recently been discussed by Paprotzki et al. [76]. Certain terms in the short-time expansion cancel out for initial states that are diagonal in the computational basis, yet play a role for, e.g., Slater determinants in the electronic quasimomentum eigenbasis. For other initial states with a finite density of electrons, e.g., a Fermi sea, phononic occupations may even become negative at large values of $\unit\omega_{0}\sim 2\unit{}$ [76]. This relates to the well-known fact that in the long-time, stationary state, occupations obtained from Ehrenfest dynamics may deviate from the correct result [77, 78].
Concerning the mere convergence with respect to its actual numerical control parameters, the statistical and numerical errors controlled by $N_{\mathrm{traj}}$ and $\Delta t$, convergence with respect to these parameters can easily be achieved in MTE. This makes MTE a technically simple-to-implement and to control technique, rendering it a reasonable starting point for hybrid methods for quantum many-body problems. While beyond the scope of paper, we envision that in the future, Lanczos or TEBD may also be combined with coupled trajectory methods like multi-configurational Ehrenfest [89, 90] which at least has internal control parameters that allow convergence to the physically correct results in principle, as was demonstrated in [97].